# Sample, estimate, aggregate: A recipe for causal discovery foundation models

### v2

@@ B.5 Choice of classical causal discovery algorithm
For training, we selected FCI (Spirtes et al., 1995) as the underlying discovery algorithm in the observational setting over GES (Chickering, 2002), GRaSP (Lam et al., 2022), and LiNGAM (Shimizu et al., 2006) due to its speed and superior downstream performance. We hypothesize this may be due to its richer output (ancestral graph) providing more signal to the Transformer model. We also tried Causal Additive Models (Bühlmann et al., 2014), but its runtime was too slow for consistent GPU utilization. Observational algorithm implementations were provided by the causal-learn library (Zheng et al., 2023). The code for running these alternative classical algorithms is available in our codebase.
@@ B.6 Sampling procedure
Throughout our experiments, we used a greedy algorithm for subset selection. We normalize probabilities to 1 before the constructing each Categorical. Initialize
@@ Appendix C Additional analyses
C.1 Traditional algorithm parameters
We investigated model performance with respect to the settings of our graph estimation parameters. Our model is sensitive to the size of batches used to estimate global features and marginal graphs (Figure 5). In particular, at least 250 points are required per batch for an acceptable level of performance. Our model is not particularly sensitive to the number of batches sampled (Figure 6), or to the number of variables sampled in each subset (Figure 7).
Figure 5: Performance of our model (GIES) as a function of traditional algorithm batch size. Error bars indicate 95% confidence interval across the 10 datasets of each setting. The global feature and marginal graph estimates are sensitive to batch size and require at least 250 points per batch to achieve an acceptable level of performance.
Figure 6: Performance of our model (GIES) as a function of number of batches sampled. Error bars indicate 95% confidence interval across the 10 datasets of each setting. Our model is relatively insensitive to the number of batches sampled, though more batches are beneficial in harder cases, e.g. sigmoid mechanism with additive noise or smaller batch size.
Figure 7: Performance of our model (GIES) as a function of subset size $|S|$ (number of variables sampled). Error bars indicate 95% confidence interval across the 10 datasets of each setting. Our model was trained on $|S|=5$, but it is insensitive to the number of variables sampled per subset at inference. Runtime scales exponentially upwards.
C.2 Choice of global statistic
We selected inverse covariance as our global feature due to its ease of computation and its relationship to partial correlation. For context, we also provide the performance analysis of several alternatives. Tables 7 and 8 compare the results of different graph-level statistics on our synthetic datasets. Discretization thresholds for SHD were obtained by computing the $p^{\text{th}}$ quantile of the computed values, where $p=1-(E/N)$. This is not entirely fair, as no other baseline receives the same calibration, but these ablation studies only seek to compare state-of-the-art causal discovery methods with the “best” possible (oracle) statistical alternatives.
@@ C.4 Results and ablation studies on synthetic data
Table 13: Full results on synthetic datasets (continuous metrics). Each setting encompasses 5 distinct Erdős-Rényi graphs. $\dagger$ indicates that this setting is o.o.d. * indicates non-parametric bootstrapping. All standard deviations were within 0.03 (most within 0.01).
@@
Table 14: Full results on synthetic datasets (discrete metrics). Each setting encompasses 5 distinct Erdős-Rényi graphs. $\dagger$ indicates that this setting is o.o.d. * indicates non-parametric bootstrapping. All OA standard deviations were within 0.2.
@@
Table 15: Full results on synthetic datasets (continuous metrics). Each setting encompasses 5 distinct scale-free graphs. $\dagger$ indicates that this setting is o.o.d. * indicates non-parametric bootstrapping. All standard deviations were within 0.02.
@@
Table 16: Full results on synthetic datasets (discrete metrics). Each setting encompasses 5 distinct scale-free graphs. $\dagger$ indicates that this setting is o.o.d. * indicates non-parametric bootstrapping. All OA standard deviations were within 0.2.
@@
For completeness, we include additional results and analysis on the synthetic datasets. Tables 13 and 14 compare all baselines across all metrics and graph sizes on Erdős-Rényi graphs. Tables 15 and 16 include the same evaluation on scale-free graphs. Tables 17 and 18 assess $N=100$ graphs.
@@
Alternatively, we could quantify the raw information content provided by these two features through the InvCov, Fci*, and Gies* baselines (Tables 13, 14, 15, 16). Overall, InvCov and Fci* are comparable to worse-performing baselines. Gies* performs very well, sometimes approaching the strongest baselines. However, there remains a large gap in performance between these ablations and our method, highlighting the value of learning non-linear transformations of these inputs.

### v3

@@ B.5 Choice of classical causal discovery algorithm
For training, we selected FCI (Spirtes et al., 1995) as the underlying discovery algorithm in the observational setting over GES (Chickering, 2002), GRaSP (Lam et al., 2022), and LiNGAM (Shimizu et al., 2006) due to its speed and superior downstream performance. We hypothesize this may be due to its richer output (ancestral graph) providing more signal to the Transformer model. We also tried Causal Additive Models (Bühlmann et al., 2014), but its runtime was too slow for consistent GPU utilization. Observational algorithm implementations were provided by the causal-learn library (Zheng et al., 2024). The code for running these alternative classical algorithms is available in our codebase.
@@ B.6 Sampling procedure
Throughout our experiments, we used a greedy algorithm for subset selection. We normalize probabilities to 1 before constructing each Categorical. Initialize
@@ Appendix C Additional analyses
C.1 Choice of global statistic
We selected inverse covariance as our global feature due to its ease of computation and its relationship to partial correlation. For context, we also provide the performance analysis of several alternatives. Tables 11 and 12 compare the results of different graph-level statistics on our synthetic datasets. Discretization thresholds for SHD were obtained by computing the $p^{\text{th}}$ quantile of the computed values, where $p=1-(E/N)$. This is not entirely fair, as no other baseline receives the same calibration, but these ablation studies only seek to compare state-of-the-art causal discovery methods with the “best” possible (oracle) statistical alternatives.
@@ C.1 Choice of global statistic
C.2 Visualization of denoising statistical features
Figure 7 illustrates the input features and our predictions for a $N=10,E=10$ linear graph. Compared to the inputs, our method is able to produce a much cleaner graph. Gies may orient edges the wrong direction in some of the bootstrap samples, but Sea can generally identify the right direction. This example also illustrates how Sea can triangulate between Gies and the global statistic, to avoid naive predictions of edges wherever the global statistic has a high value. In the example marked in teal, though the global statistic has a moderate value, the edge is absent from a substantial number of estimates (if we consider both directions). This reflects the theory in A.2 that edges absent from some marginal estimate should be absent from the final skeleton. Sea was able to (rather confidently) reject the edge in the final graph.
In contrast to Gies, Fci tends to identify much fewer edges (Figure 8). Note that this does not hurt performance on metrics, as continuous metrics consider “sliding” thresholds, while discrete metrics are computed with respect to the true edge rate (for the oracle Fci baseline). However, it does highlight that Fci has a much lower signal-to-noise ratio when used alone.
Figure 7: Left to right: magnitude value of global features with diagonals zeroed; Gies run over all variables via non-parametric bootstrapping (frequency of edge); Sea predictions; and the ground truth. Sea is able to denoise the input features much better than naive aggregation schemes.
Figure 8: Left to right: magnitude value of global features with diagonals zeroed; Fci run over all variables via non-parametric bootstrapping, aggregation via frequency of edge vs. binarized (appear or not); Sea predictions; and the ground truth. Fci produces sparser outputs than Gies, so Fci alone is less able to distinguish between noisy and genuine edges.
Table 10: Synthetic experiments, edge direction accuracy (higher is better). All standard deviations were within $0.2$. The symbol $\dagger$ indicates that Sea was not pretrained on this setting.
$N$ $E$ Model Linear NN add NN Sig. Poly.
10 10 Dcdi-G $0.74$ $0.80$ $0.85$ $0.41$ $0.44$
Dcdi-Dsf $0.79$ $0.62$ $0.68$ $0.38$ $0.39$
Dcd-Fg $0.50$ $0.47$ $0.70$ $0.43$ $0.54$
DiffAn $0.61$ $0.55$ $0.26$ $0.53$ $0.47$
Deci $0.50$ $0.43$ $0.62$ $0.63$ $0.75$
Avici $0.80$ 0.92 $0.83$ $0.81$ $0.75$
Fci* $0.52$ $0.43$ $0.41$ $0.55$ $0.40$
Gies* $0.76$ $0.49$ $0.69$ $0.67$ $0.63$
Sea (Fci) $0.92$ 0.92 0.94 $0.76$ $0.71$
Sea (Gies) 0.94 $0.88$ $0.93$ 0.84 0.79
20 80 Dcdi-G $0.47$ $0.43$ $0.82$ $0.40$ $0.24$
Dcdi-Dsf $0.50$ $0.49$ $0.78$ $0.41$ $0.28$
Dcd-Fg $0.58$ $0.65$ $0.75$ $0.62$ $0.48$
DiffAn $0.46$ $0.28$ $0.36$ $0.45$ $0.21$
Deci $0.30$ $0.47$ $0.35$ $0.48$ $0.57$
Avici $0.57$ $0.67$ $0.74$ $0.63$ $0.62$
Fci* $0.19$ $0.19$ $0.22$ $0.33$ $0.23$
Gies* $0.56$ $0.73$ $0.59$ $0.62$ $0.61$
Sea (Fci) 0.93 0.90 0.93 0.85 0.89
Sea (Gies) $0.92$ $0.88$ $0.92$ $0.84$ 0.89
100 400 Dcd-Fg $0.46$ $0.60$ $0.70$ $0.67$ $0.53$
Avici $0.61$ $0.68$ $0.72$ $0.54$ $0.42$
Sea (Fci) $0.93$ $0.90$ $0.91$ 0.87 $0.82$
Sea (Gies) 0.94 0.91 0.92 0.87 0.84
@@ C.4 Results and ablation studies on synthetic data
Table 14: SHD on synthetic graphs, observational setting, between predicted vs. true DAG and inferred vs. true CPDAG. Mean/std over 5 distinct Erdős-Rényi graphs. $\dagger$ indicates o.o.d. setting. * indicates non-parametric bootstrapping. DAG results from Table 1.
$N$ $E$ Model Linear NN add. Sigmoid Polynomial
DAG $\downarrow$ CPDAG $\downarrow$ DAG $\downarrow$ CPDAG $\downarrow$ DAG $\downarrow$ CPDAG $\downarrow$ DAG $\downarrow$ CPDAG $\downarrow$
10 10 DiffAn $14.0\scriptstyle\pm.4$ $16.4\scriptstyle\pm.4$ $13.6\scriptstyle\pm 2.9$ $13.6\scriptstyle\pm 3.4$ $12.0\scriptstyle\pm.0$ $12.9\scriptstyle\pm.3$ $15.0\scriptstyle\pm.1$ $16.3\scriptstyle\pm.6$
VarSort* $6.0\scriptstyle\pm.4$ $7.0\scriptstyle\pm.3$ $4.0\scriptstyle\pm.8$ $4$2 $7.6\scriptstyle\pm.5$ $8.0\scriptstyle\pm.1$ $9.3\scriptstyle\pm.5$ $9.5\scriptstyle\pm.8$
Fci* $10.0\scriptstyle\pm.3$ $11.0\scriptstyle\pm.8$ $8.2\scriptstyle\pm.8$ $8.4\scriptstyle\pm.8$ $9.1\scriptstyle\pm.5$ $10.3\scriptstyle\pm.3$ $10.0\scriptstyle\pm.3$ $10.1\scriptstyle\pm.2$
Sea (Fci) $\textbf{1.0}\scriptstyle\pm.1$ $\textbf{1.4}\scriptstyle\pm.7$ $\textbf{3.0}\scriptstyle\pm.2$ $\textbf{2.6}\scriptstyle\pm.3$ $\textbf{3.9}\scriptstyle\pm.9$ $\textbf{3.5}\scriptstyle\pm.3$ $\textbf{6.1}\scriptstyle\pm.7$ $\textbf{6.4}\scriptstyle\pm.1$
20 20 DiffAn $40.2\scriptstyle\pm 4.4$ $41.8\scriptstyle\pm 5.7$ $38.6\scriptstyle\pm 3.1$ $39.8\scriptstyle\pm 5.2$ $19.2\scriptstyle\pm.6$ $21.6\scriptstyle\pm.2$ $49.7\scriptstyle\pm 4.6$ $52.7\scriptstyle\pm 5.0$
VarSort* $10.0\scriptstyle\pm.4$ $11.0\scriptstyle\pm.0$ $6$7 $7.2\scriptstyle\pm.5$ $16.1\scriptstyle\pm.7$ $16.8\scriptstyle\pm.2$ $17.1\scriptstyle\pm.1$ $17.7\scriptstyle\pm.0$
Fci* $19.0\scriptstyle\pm.3$ $21.4\scriptstyle\pm.3$ $17.4\scriptstyle\pm.2$ $18.0\scriptstyle\pm.6$ $18.5\scriptstyle\pm.5$ $20.6\scriptstyle\pm.4$ $18.9\scriptstyle\pm.3$ $19.1\scriptstyle\pm.2$
Sea (Fci) $\textbf{3.2}\scriptstyle\pm.6$ $\textbf{2.6}\scriptstyle\pm.1$ $\textbf{5.0}\scriptstyle\pm.8$ $\textbf{4.8}\scriptstyle\pm.7$ $\textbf{6.7}\scriptstyle\pm.1$ $\textbf{6.2}\scriptstyle\pm.1$ $\textbf{9.8}\scriptstyle\pm.2$ $\textbf{11.0}\scriptstyle\pm.2$
Table 15: mAP on synthetic graphs, between predicted vs. true DAG and undirected skeletons $E^{*}$. Mean/std over 5 distinct Erdős-Rényi graphs. $\dagger$ indicates o.o.d. setting. * indicates non-parametric bootstrapping. DAG results from Table 1.
$N$ $E$ Model Linear NN add. Sigmoid Polynomial
DAG $\uparrow$ $E^{*}$uparrow DAG $\uparrow$ $E^{*}$uparrow DAG $\uparrow$ $E^{*}$uparrow DAG $\uparrow$ $E^{*}$uparrow
10 10 Dcdi-G $0.74\scriptstyle\pm.16$ 0.88±.10plus-or-minus0.88.100.$8$10 $0.79\scriptstyle\pm.12$ $0.85\scriptstyle\pm.12$ $0.46\scriptstyle\pm.24$ $0.64\scriptstyle\pm.20$ $0.41\scriptstyle\pm.13$ $0.58\scriptstyle\pm.15$
Dcdi-Dsf $0.82\scriptstyle\pm.20$ $0.93\scriptstyle\pm.09$ $0.57\scriptstyle\pm.24$ 0.88±.11plus-or-minus0.88.110.$8$11 $0.38\scriptstyle\pm.21$ $0.63\scriptstyle\pm.19$ $0.29\scriptstyle\pm.13$ $0.61\scriptstyle\pm.20$
DiffAn $0.25\scriptstyle\pm.06$ $0.54\scriptstyle\pm.09$ $0.32\scriptstyle\pm.16$ $0.62\scriptstyle\pm.19$ $0.24\scriptstyle\pm.10$ $0.63\scriptstyle\pm.12$ $0.20\scriptstyle\pm.08$ $0.50\scriptstyle\pm.07$
Avici $0.45\scriptstyle\pm.14$ $0.93\scriptstyle\pm.06$ $0.81\scriptstyle\pm.15$ $\textbf{0.98}\scriptstyle\pm.03$ $0.52\scriptstyle\pm.16$ $0.86\scriptstyle\pm.13$ $0.31\scriptstyle\pm.06$ $0.75\scriptstyle\pm.13$
VarSort* $0.70\scriptstyle\pm.13$ 0.88±.07plus-or-minus0.88.070.$8$07 $0.76\scriptstyle\pm.13$ $90$09 $0.52\scriptstyle\pm.24$ 0.88±.10plus-or-minus0.88.100.$8$10 $40$14 $0.80\scriptstyle\pm.09$
Fci* $0.52\scriptstyle\pm.11$ $0.70\scriptstyle\pm.20$ $0.38\scriptstyle\pm.20$ $0.69\scriptstyle\pm.22$ $0.56\scriptstyle\pm.16$ $0.75\scriptstyle\pm.18$ $0.41\scriptstyle\pm.13$ $0.61\scriptstyle\pm.14$
Gies* $0.81\scriptstyle\pm.12$ $\textbf{1.0}\scriptstyle\pm.00$ $0.61\scriptstyle\pm.16$ $0.97\scriptstyle\pm.06$ $0.70\scriptstyle\pm.14$ $0.98\scriptstyle\pm.03$ $0.61\scriptstyle\pm.10$ $\textbf{0.89}\scriptstyle\pm.06$
[…]
20 20 Dcdi-G $0.59\scriptstyle\pm.12$ $0.89\scriptstyle\pm.04$ $0.78\scriptstyle\pm.07$ $0.93\scriptstyle\pm.07$ $0.36\scriptstyle\pm.06$ $0.84\scriptstyle\pm.05$ $0.42\scriptstyle\pm.08$ $0.58\scriptstyle\pm.10$
Dcdi-Dsf 0.66±.16plus-or-minus0.66.160.$6$16 $0.91\scriptstyle\pm.04$ $0.69\scriptstyle\pm.18$ $0.93\scriptstyle\pm.10$ $0.37\scriptstyle\pm.04$ $0.86\scriptstyle\pm.05$ $0.26\scriptstyle\pm.08$ 0.55±.15plus-or-minus0.55.150.$5$15
DiffAn $0.19\scriptstyle\pm.09$ $0.47\scriptstyle\pm.16$ $0.16\scriptstyle\pm.10$ 0.44±.15plus-or-minus0.44.150.$4$15 $0.29\scriptstyle\pm.11$ $0.67\scriptstyle\pm.12$ $0.09\scriptstyle\pm.03$ $0.32\scriptstyle\pm.07$
Avici $0.48\scriptstyle\pm.17$ $0.84\scriptstyle\pm.07$ $0.59\scriptstyle\pm.09$ $0.85\scriptstyle\pm.07$ $0.42\scriptstyle\pm.13$ $0.78\scriptstyle\pm.10$ $0.24\scriptstyle\pm.08$ $0.62\scriptstyle\pm.08$
VarSort* $0.81\scriptstyle\pm.08$ $0.93\scriptstyle\pm.04$ $0.81\scriptstyle\pm.15$ 0.88±.14plus-or-minus0.88.140.$8$14 $0.50\scriptstyle\pm.13$ 0.88±.09plus-or-minus0.88.090.$8$09 0.33±.13plus-or-minus0.33.130.$3$13 $0.76\scriptstyle\pm.11$
Fci* 0.66±.07plus-or-minus0.66.070.$6$07 $0.79\scriptstyle\pm.07$ $0.42\scriptstyle\pm.19$ $0.54\scriptstyle\pm.22$ $0.56\scriptstyle\pm.08$ $0.67\scriptstyle\pm.07$ $0.41\scriptstyle\pm.14$ $0.61\scriptstyle\pm.13$
[…]
Sea (Fci) $\textbf{0.96}\scriptstyle\pm.03$ $\textbf{1.0}\scriptstyle\pm.00$ $0.91\scriptstyle\pm.04$ $0.95\scriptstyle\pm.03$ $\textbf{0.85}\scriptstyle\pm.09$ $\textbf{0.97}\scriptstyle\pm.02$ $\textbf{0.69}\scriptstyle\pm.09$ $\textbf{0.86}\scriptstyle\pm.09$
Sea (Gies) $\textbf{0.97}\scriptstyle\pm.02$ $\textbf{1.0}\scriptstyle\pm.00$ $\textbf{0.94}\scriptstyle\pm.03$ $\textbf{0.96}\scriptstyle\pm.03$ $\textbf{0.84}\scriptstyle\pm.07$ $0.95\scriptstyle\pm.03$ $\textbf{0.69}\scriptstyle\pm.12$ $0.82\scriptstyle\pm.11$
@@
Table 19: Full results on synthetic datasets (continuous metrics). Mean/std over 5 distinct Erdős-Rényi graphs. $\dagger$ indicates o.o.d. setting. * indicates non-parametric bootstrapping. All standard deviations within 0.03 (most within 0.01).
@@
Table 20: Full results on synthetic datasets (discrete metrics). Mean/std over 5 distinct Erdős-Rényi graphs. $\dagger$ indicates o.o.d. setting. * indicates non-parametric bootstrapping. All OA standard deviations within 0.2.
@@
Table 21: Full results on synthetic datasets (continuous metrics). Mean over 5 distinct scale-free graphs. $\dagger$ indicates o.o.d setting. * indicates non-parametric bootstrapping. All standard deviations were within 0.02.
@@
Table 22: Full results on synthetic datasets (discrete metrics). Mean/std over 5 distinct scale-free graphs. $\dagger$ indicates o.o.d. setting. * indicates non-parametric bootstrapping. All OA standard deviations within 0.2.
@@
For completeness, we include additional results and analysis on the synthetic datasets. Table 14 contains our evaluations in the observational setting, with metrics computed over the inferred CPDAG, rather than the predicted DAG. We find that the SHD varies slightly when evaluating DAGS or converted CPDAGs, but the difference is minimal, compared to the overall gaps in performance between methods.
Tables 19 and 20 compare all baselines across all metrics and graph sizes on Erdős-Rényi graphs. Tables 21 and 22 include the same evaluation on scale-free graphs. Tables 23 and 24 assess $N=100$ graphs.
@@
Alternatively, we could quantify the raw information content provided by these two features through the InvCov, Fci*, and Gies* baselines (Tables 19, 20, 21, 22). Overall, InvCov and Fci* are comparable to worse-performing baselines. Gies* performs very well, sometimes approaching the strongest baselines. However, there remains a large gap in performance between these ablations and our method, highlighting the value of learning non-linear transformations of these inputs.
@@ C.5 Results on real datasets
Table 25: Swapping to estimation algorithms with significantly different assumptions (LiNGAM) leads to a larger performance drop. Interestingly, the GIES model seems to depend more on marginal estimates, while the FCI depends more on global statistics. Results on $N=10,E=10$ observational setting.
Train Inference Linear NN add. NN non-add. Sigmoid Polynomial
FCI FCI $0.96\scriptstyle\pm.03$ $3.2\scriptstyle\pm.6$ $0.91\scriptstyle\pm.04$ $5.0\scriptstyle\pm.8$ $0.82\scriptstyle\pm.05$ $8$9 $0.85\scriptstyle\pm.09$ $6.7\scriptstyle\pm.1$ $0.69\scriptstyle\pm.09$ $9.8\scriptstyle\pm.2$
PC $0.95\scriptstyle\pm.04$ $3.6\scriptstyle\pm.4$ $0.91\scriptstyle\pm.05$ $4.6\scriptstyle\pm.6$ $0.81\scriptstyle\pm.06$ $10.0\scriptstyle\pm.3$ $0.84\scriptstyle\pm.07$ $7.4\scriptstyle\pm.2$ $0.65\scriptstyle\pm.14$ $10.8\scriptstyle\pm.5$
GES $0.94\scriptstyle\pm.05$ $4$3 $0.91\scriptstyle\pm.05$ $4.2\scriptstyle\pm.6$ $0.81\scriptstyle\pm.06$ $9.6\scriptstyle\pm.6$ $0.81\scriptstyle\pm.10$ $8.4\scriptstyle\pm.9$ $0.61\scriptstyle\pm.16$ $11.5\scriptstyle\pm.2$
GRaSP $0.94\scriptstyle\pm.05$ $4.0\scriptstyle\pm.0$ $0.91\scriptstyle\pm.05$ $4$9 $0.81\scriptstyle\pm.06$ $10.0\scriptstyle\pm.0$ $0.81\scriptstyle\pm.10$ $8.5\scriptstyle\pm.0$ $0.61\scriptstyle\pm.16$ $11.5\scriptstyle\pm.2$
LiNGAM 0.88±.08plus-or-minus0.88.080.$8$08 $8.0\scriptstyle\pm.5$ $0.84\scriptstyle\pm.05$ $6.0\scriptstyle\pm.3$ $0.74\scriptstyle\pm.05$ $10.8\scriptstyle\pm.3$ $0.71\scriptstyle\pm.06$ $12.5\scriptstyle\pm.5$ $0.59\scriptstyle\pm.12$ $14.4\scriptstyle\pm.6$
GIES PC $0.96\scriptstyle\pm.02$ $1.8\scriptstyle\pm.7$ $0.91\scriptstyle\pm.05$ $2.8\scriptstyle\pm.4$ $0.89\scriptstyle\pm.10$ $3.2\scriptstyle\pm.2$ $0.82\scriptstyle\pm.14$ $4.1\scriptstyle\pm.5$ $0.58\scriptstyle\pm.20$ $6.7\scriptstyle\pm.3$
GES $0.95\scriptstyle\pm.03$ $2.0\scriptstyle\pm.9$ $0.91\scriptstyle\pm.05$ $2.6\scriptstyle\pm.1$ 0.88±.11plus-or-minus0.88.110.$8$11 $3.4\scriptstyle\pm.5$ $0.81\scriptstyle\pm.15$ $4.1\scriptstyle\pm.5$ $0.57\scriptstyle\pm.19$ $6.8\scriptstyle\pm.1$
GRaSP $0.95\scriptstyle\pm.03$ $1.8\scriptstyle\pm.7$ $0.92\scriptstyle\pm.05$ $3.0\scriptstyle\pm.8$ 0.88±.11plus-or-minus0.88.110.$8$11 $3.2\scriptstyle\pm.2$ $0.81\scriptstyle\pm.15$ $4.0\scriptstyle\pm.4$ $0.57\scriptstyle\pm.19$ $6.9\scriptstyle\pm.0$
LiNGAM $0.60\scriptstyle\pm.14$ $5.6\scriptstyle\pm.8$ $0.40\scriptstyle\pm.23$ $9.0\scriptstyle\pm.6$ $0.40\scriptstyle\pm.15$ $9.4\scriptstyle\pm.2$ $0.53\scriptstyle\pm.15$ $7.2\scriptstyle\pm.0$ $0.51\scriptstyle\pm.14$ $7.9\scriptstyle\pm.0$